# OTTER: Improving Zero-Shot Classification via Optimal Transport

### v1

@@ D.2 Zero-shot classification setup
Table 12: Accuracy in the prompt-enhanced zero-shot classification by Classification by Description (CD) [38]. We can observe OTTER’s capability to provide further enhancements upon refined the improvements achieved through refined prompts.
D.4 Detailed experiment results of Section 5.2.2
Class balance estimation errors
We report the class balance errors in Section 5.2.2. As a metric, we use the total variance $TV(\nu,\hat{\nu})=\frac{1}{2}\left|\left|\nu-\hat{\nu}\right|\right|_{1}$. We use zeroshot prediction scores and linear probing prediction scores for BBSE. $\hat{\nu}^{zs}$ denotes the estimated class balance based on zero-shot prediction scores, and $\hat{\nu}^{lp}$ represents the estimated class balance based on linear probing prediction scores.
@@
Table 13: Class balance estimation error with BBSE in Section 5.2.2. We report the mean of 10 different random samplings of the validation set. Lower is better.
@@
We provide full results of Section 5.2.2. Specifically, we additionally report the results of combination with linear probing in text classification tasks and the results of zero-shot classification results in image classification tasks.
@@
Dataset ZS ZS BBSE+PM ZS BBSE+OT LP LP BBSE+PM LP BBSE+OT
CIFAR10 88.3 72.7 87.5 90.2 89.8 90.0
CIFAR100 63.8 3.2 59.1 58.3 24.4 60.5
Caltech101 79.8 32.5 80.7 91.5 87.5 91.4
Caltech256 79.8 6.0 80.3 84.5 58.4 85.4
Country211 19.8 1.5 15.9 12.4 9.2 13.2
DTD 39.0 3.2 31.2 58.6 49.0 59.3
EUROSAT 32.9 19.2 34.0 74.6 71.6 75.9
Flowers102 64.0 40.3 60.8 89.0 87.8 90.2
Food101 85.6 15.3 82.3 79.1 60.6 79.8
Oxford-IIIT-Pet 83.8 43.3 71.4 75.7 72.0 75.6
Stanford-Cars 55.7 2.3 51.7 64.5 65.4 66.3
STL10 98.0 97.4 96.9 97.7 97.5 97.6
SUN397 47.1 6.9 25.6 0.2 0.2 0.2
cub 46.0 3.3 45.5 72.2 63.3 75.6
ImageNet 60.2 0.8 57.7 56.8 53.6 59.8
ImageNet-r 68.9 1.7 63.3 54.9 47.6 57.1
ImageNet-Sketch 39.8 0.8 40.4 43.4 37.9 48.3
Amazon 74.0 47.9 89.1 71.3 66.9 71.3
CivilComments 48.3 69.1 55.8 53.8 45.5 53.8
Gender 84.0 57.0 87.8 78.0 71.2 78.5
HateXplain 30.4 34.4 35.2 32.8 32.7 32.3
Table 14: Accuracy (%) with OTTER combined with class balance estimation. ZS BBSE denotes BBSE class balance estimation based on zero-shot prediction scores, and LP BBSE denotes BBSE class balance estimation based on linear probing prediction scores. We report the mean of 10 different random samplings of the validation set.
@@
The results, as depicted in Figure 5, corroborate our hypothesis. It is evident that the error in class balance estimation diminishes with an increasing number of samples, leading to a concurrent enhancement in the accuracy of OTTER.
@@
D.5 Detailed experiment setup and results of Section 5.2.3
@@
We report the class balance estimation error in Section 5.2.3. Table 15 shows the total variation between true class balance and estimated class balance. We can expect a significant accuracy improvement for RN50, RN101, and ViT-B/16 based on this table.
@@
Table 15: Class balance estimation error in the Section 5.2.3 experiment. Class balance estimation error is measured by total variation distance. We report the mean of 10 different samplings of the validation set. H-BBSE denotes the class balance estimation using hierarchy upon BBSE.

### v2

@@ D.2 Zero-shot classification setup
Table 12: Accuracy in the prompt-enhanced zero-shot classification by Classification by Description (CD) [menon2022visual]. We can observe OTTER’s capability to provide further enhancements upon refined the improvements achieved through refined prompts.
D.4 Detailed experiment results of Section 5.3
Label distribution estimation errors
We report the label distribution estimation errors in Section 5.3. As a metric, we use the total variance $TV(\nu,\hat{\nu})=\frac{1}{2}\left|\left|\nu-\hat{\nu}\right|\right|_{1}$. We use zeroshot prediction scores and linear probing prediction scores for BBSE. $\hat{\nu}^{zs}$ denotes the estimated class balance based on zero-shot prediction scores, and $\hat{\nu}^{lp}$ represents the estimated class balance based on linear probing prediction scores.
@@
Table 13: Class balance estimation error with BBSE in Section 5.3. We report the mean of 10 different random samplings of the validation set. Lower is better.
@@
We provide full results of Section 5.3. Specifically, we additionally report the results of combination with linear probing in text classification tasks and the results of zero-shot classification results in image classification tasks.
@@
The results, as depicted in Figure 5, corroborate our hypothesis. It is evident that the error in class balance estimation diminishes with an increasing number of samples, leading to a subsequent enhancement in the accuracy of OTTER.
@@
D.5 Detailed experiment setup and results of Section 5.4
@@
We report the class balance estimation error in Section 5.4. Table 14 shows the total variation between true class balance and estimated class balance. We can expect a significant accuracy improvement for RN50, RN101, and ViT-B/16 based on this table.
@@
Table 14: Class balance estimation error in the Section 5.4 experiment. Class balance estimation error is measured by total variation distance. We report the mean of 10 different samplings of the validation set. H-BBSE denotes the class balance estimation using hierarchy upon BBSE.
[…]
Setup and Procedure.
We use our synthetic experiment setup (Section 5.2) with perturbation noise $\delta=0$ and label distribution $\alpha=0$. Additionally, we generated a validation set that follows the same distribution as the test set. After learning the reweighting parameter $r$ in the validation set using $y_{otter}$ as pseudolabels, we evaluated R-OTTER on the test set, comparing the results to those of zero-shot and OTTER. We expect that, if successful, R-OTTER will similarly gain improvement over zero-shot when the source and target distributions increasingly differ.
Results
[…]